# A Matrix Model Proposal for Quantum Gravity and the Quantum Mechanics of Black Holes

### v2

@@ Abstract
We propose a quantum mechanical theory of quantum spaces described by large $N$ noncommutative geometry as a model for quantum gravity. The theory admits fuzzy sphere as static solution. Over the fuzzy geometry, the quantum mechanics of the fermions is given by a sum of oscillators with equal frequency. The energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy. This state of the fuzzy sphere obeys the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere is identified with the black hole horizon. Moreover the set of states in the Fermi sea gives precisely the Bekenstein-Hawking entropy. We thus propose that quantum black holes are described by fuzzy spheres with a half-filled Fermi sea in our theory. We also consider a system of two fuzzy spheres by embedding them as blocks in the matrix quantum mechanics. When the distance $r$ between the two fuzzy spheres is small, the total energy of the system can be computed using perturbation theory. We show that in the leading order of large $N$ limit, the interaction energy depends on the product of the black holes masses and the Newton constant exactly as in Newton gravity. To extract the correct $r$ dependence in the long range, a resummation of all the large $N$ corrections is needed. We outline how Newton gravity for static sources may be reproduced in the long distance limit. We also show that the interaction energy is generally finite in the short distance limit, suggesting that the black hole singularity in general relativity is removed in our quantum description of black hole.
@@ 1 Introduction
where $h(X)$ denotes some Yukawa coupling and $V(X)$ the self-interaction. Our goal was to identify basic properties of quantum gravity which are essential to the construction of the theory. Our hope was that the model would be helpful in ways similar to that of the Bohr atomic model to the building of quantum mechanics. We found [19] that if our model admits a constant density of energy eigenstates and if the Fermi sea of the system is filled up to a Fermi energy level that is inversely proportional to the system size, then the Schwarzschild radius of black hole is reproduced for the system. Moreover the system is in a highly degenerate energy state whose counting of microstates gives precisely the Bekenstein-Hawking entropy. While the result is quite encouraging, the identified conditions are sufficient ones, and it is not clear if all the assumptions can indeed be implemented consistently in a quantum mechanics. The construction of a consistent quantum mechanical theory of spacetime which includes quantum black hole is the main motivation of this work.
@@
In this work, we propose an explicit large $N$ quantum mechanics of non-abelian bosonic and fermionic variables as a model of quantum gravity. The use of large $N$ is inspired by the previous success of the BFSS and IKKT matrix models. There are three bosonic variables $X^{a}$ in the adjoint of $SU(N)$ which correspond to the three spatial coordinates of a quantized noncommutative space. We note that the emergence of noncommutative geometry is generally expected when space (or spacetime) is quantized in quantum gravity  [^1] The studied of quantized spacetime in terms of noncommutative geometry goes back to the study of Snyder [21] and Yang [22]. . In addition, there are two fermionic variables which couple minimally to the geometry, and make up the fundamental Hilbert space of quantum gravity. We show that our theory of quantized spacetime admits fuzzy sphere as static solution. Over the fuzzy geometry, the quantum mechanics of the fermions gives a Fermi sea with uniform energy levels. It is amazing that the energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy and satisfies the the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere is identified with the black hole horizon. Moreover the set of states in the half-filled Fermi sea gives precisely the Bekenstein-Hawking entropy. We therefore propose that quantum black holes are described by fuzzy spheres with a half-filled Fermi sea in our theory. Quite amazingly, the quantum mechanics proposed here realizes pretty much all the ideas outlined earlier in [18, 19, 20].
@@
To show that the fuzzy spheres are gravitating objects, we need to show that the interaction between them agrees with gravity, e.g. with Newton gravity in the long distance limit. We thus consider a system of two fuzzy spheres by embedding them as blocks in the matrix quantum mechanics. When the distance $r$ between the two fuzzy sphere is small, the total energy of the system can be computed using perturbation theory. We show that in the leading order of large $N$ limit, the interaction energy depends on the product of black holes masses and the Newton constant exactly as is expected in Newton gravity. We also discuss how a resummation of all the large $N$ corrections could reproduce the correct Newton limit at large distance. It is interesting to note that the interaction energy is generally finite in the short distance limit, suggesting that the black hole singularity in general relativity is removed in our quantum description of black hole.
@@
The plan of the paper is as follows. In section 2, we give our proposal of a large $N$ quantum mechanics as a fundamental formulation of quantum gravity in 3-dimensions. In section 3, we show that the fuzzy sphere solution with a half-filled Fermi sea has the desired properties of a quantum black hole, namely with the black hole mass-radius relation and the Bekenstein-Hawking black hole entropy reproduced correctly. In section 4, we consider a system of 2 fuzzy spheres and derive their interaction energy in the small separation limit. We discuss how a large $N$ resummation of all the higher order perturbations including the off-diagonal block interaction could generate the desired gravitational forces between the quantum black holes. Discussion is found in section 5.
@@ 2 A proposed model of quantum gravity
Let us consider a theory of quantized space where the spatial coordinates becomes operators $X^{a}$ ($a=1,2,3$) and are represented by $SU(N)$ matrices. Here $N$ is large and characterizes the dimensions of the Hilbert space of quantum gravity. In addition, we propose to include fermionic degrees of freedom in the fundamental formulation of quantum gravity. In our model proposed below, they are given by the 2-components spinors $\psi^{\dagger},\psi$ in the adjoint representation of $SU(N)$. Our choice of degrees of freedom is a minimalistic one and is allowed since we do not assume supersymmetry. In fact, it is important to realize that going bottom-up, there is no physical reason to insist on having supersymmetry. We remark that in the usual top down approach, supersymmetry is often needed for the reason of consistency of formulation. e.g. for string/M theory in 10 or 11 dimensions. However, supersymmetry is irrelevant for the well-definedness and consistency of quantum mechanics as there is no issue of renormalizability here. Also the usual advantages of supersymmetric quantum mechanics such as exact computation of quantities such as index or partition function does not appeal to us since here we are after the construction of a physical theory rather than a theory that is exact or calculable. Therefore we will not insist on having supersymmetry. This allows for a much wider choice of terms in the Lagrangian which are forbidden otherwise. We propose the following definition for the quantum theory of spacetime,
$L=\mbox{tr}\left(\frac{1}{2M_{0}}\dot{X}^{a2}+\frac{M_{P}}{N^{2}}[X^{a},X^{b}]^{2}+\frac{4M_{P}}{N^{2}}X^{a2}+i\dot{\psi}^{\dagger}\psi-\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi-M_{P}{\bf 1}\right).$ (2.1)
Here $X^{a}_{mn},\psi_{mn},\psi_{mn}^{\dagger}$ are dimensionless, $\sigma^{a}$ are the Pauli matrices, $M_{0}$ and $M_{P}$ are some mass scales with $M_{P}$ being the Planck mass that is related to the Newtonian constant $G$ as
$M_{P}^{2}=\frac{2}{\pi G}.$ (2.2)
The Lagrangian includes, apart from the kinetic terms for the bosons and fermions, a Yang-Mills type potential and a Yukawa coupling. These are quite standard. In addition, there is also included a quadratic mass term. This term is forbidden by supersymmetry. In addition, we include a cosmological constant-like term. The bosonic potential of the theory is Higgs like $-\phi^{2}+\phi^{4}$. In addition to the trivial solution at the origin. There are nontrivial fuzzy spheres as nontrivial noncommutative space solutions to the theory. Due to their physical properties, we propose them as quantum description of Schwarzschild black holes.
The Lagrangian (2.1) is invariant under the global $SU(N)$ transformation
@@
The time derivative should be promoted to a covariant derivative if the $SU(N)$ symmetry becomes a time-dependent one. The theory also has an $SO(3)$ rotational invariance where $X^{i}$ transforms as a vector and $\psi$ in the spinor representation. The quantum mechanical Hamiltonian is given by
$H=\mbox{tr}\left(\frac{M_{0}}{2}P^{i2}-\frac{M_{P}}{N^{2}}[X^{a},X^{b}]^{2}-\frac{4M_{P}}{N^{2}}X^{a2}+\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi+M_{P}{\bf 1}\right).$ (2.4)
where $P^{a}=\dot{X}^{a}/M_{0}$ is the conjugate momentum for the bosonic degrees of freedom and satisfies the commutation relation
$[X^{a}_{mk},P^{b}_{nl}]=i\delta^{ab}\delta_{mn}\delta_{kl}.$ (2.5)
As for the fermionic degrees of freedom, the conjugate momentum $\pi=i\psi^{\dagger}$ give rises to the anticommutation relation
$\{\psi^{\dagger}_{mk\alpha},\psi_{nl\beta}\}=\delta_{mn}\delta_{kl}\delta_{\alpha\beta},\quad\alpha,\beta=1,2.$ (2.6)
We note that in the large $N$ limit, our model is equivalent to some non-supersymmetric version of the IKKT-like instantonic model in 4 dimensions by a large $N$ reduction [23].
We propose that the large $N$ quantum mechanics (2.4) provides a fundamental formulation of quantum gravity in 3 space dimensions. We remark that our proposal of a large $N$ matrix quantum mechanics as a fundamental theory of quantum gravity is similar in philosophy to that of the M(atrix) theory [24]. However we do not assume supersymmetry, and so we can write down our model directly for 3 space dimensions, and we can include a mass term. We note that a mass term has been considered in the IKKT matrix model where interesting classical solutions such as fuzzy spacetime [25] and expanding universe [26, 27, 28, 29], have been obtained. However, our proposal of a large $N$ quantum mechanics as a fundamental formulation of quantum gravity and the connection with black hole are new.
3 Quantum black hole as fuzzy sphere with half-filled Fermi sea
@@
$-\frac{1}{M_{0}}\ddot{X}^{a}+\frac{4M_{P}}{N^{2}}\left([X^{b},[X^{a},X^{b}]]+2X^{a}\right)=0.$ (3.1)
@@
$[[X^{a},X^{b}],X^{b}]=2X^{a}.$ (3.2)
This can be solved by the spin $j=(N-1)/2$ representation of $SU(2)$, which are given by $2j+1=N$ dimensional matrices satisfying
$[X^{a},X^{b}]=i\epsilon^{abc}X^{c}.$ (3.3)
@@
the configuration (3.3), (3.4) define a fuzzy sphere. Let us introduce dimensional coordinates $Y^{a}=2l_{P}X^{a}$ for some length scale $l_{P}$, the solution becomes
$[Y^{a},Y^{b}]=\frac{2iR}{\sqrt{N^{2}-1}}\epsilon_{abc}Y^{c},\quad\sum_{a}Y^{a2}=R^{2}{\bf 1},$ (3.5)
where $R^{2}=(N^{2}-1)l_{P}^{2}$. This describes a fuzzy sphere of radius $R$. In the following, we will take $l_{P}$ to be the Planck length given by
$l_{P}=\frac{2}{\pi M_{P}}=\sqrt{\frac{2G}{\pi}}.$ (3.6)
We are interested in the energy of the fuzzy sphere system. Over the geometry of the fuzzy sphere, the Hamiltonian of the system is given by
$H=H_{B}+H_{F},$ (3.7)
where
$H_{B}=\frac{NM_{P}}{2}(1+\frac{1}{N^{2}})$ (3.8)
obtained from the constant, $X^{2}$ and $X^{4}$ terms in (2.4), is the energy arises from the fuzzy sphere geometry, and
$H_{F}=\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi$ (3.9)
is the fermionic part of the Hamiltonian. It is easy to see that $H_{F}$ describes a collection of fermionic oscillators with equal frequency. To show this, we note that for the fuzzy sphere, the matrix $K:=\sigma^{a}X^{a}$ satisfies $K^{2}-K-\frac{N^{2}-1}{4}{\bf 1}=0$ and so it has eigenvalues $(N+1)/2$ or $-(N-1)/2$. Since it is traceless, therefore $N-1$ of the eigenvalues are positive and $N+1$ of them are the negative ones. In the leading order of large $N$ limit, $K$ has the eigendecomposition
@@
Here ${\cal U}^{p}_{n\beta},{\cal V}^{p}_{n\beta},p=1,\cdots,N$ are eigenvectors of $K$ with positive and negative eigenvalue. Let us from now on consider the leading large $N$ limit. Introduce the fermionic oscillators
$\xi^{p}_{k}:={\cal U}^{p\dagger}_{n\beta}\psi_{nk\beta},\quad\chi^{p\dagger}_{k}:={\cal V}^{p\dagger}_{n\beta}\psi_{nk\beta},$ (3.11)
which satisfy
$\{\xi^{p}_{k},\xi^{q\dagger}_{l}\}=\delta^{pq}\delta_{kl},\quad\{\chi^{p}_{k},\chi^{q\dagger}_{l}\}=\delta^{pq}\delta_{kl},$ (3.12)
then $H_{F}$ takes on the form as claimed above:
$H_{F}=\frac{M_{P}}{2N}\left(\sum_{p,k=1}^{N}\xi_{k}^{p\dagger}\xi_{k}^{p}+\chi_{k}^{p\dagger}\chi_{k}^{p}-N^{2}\right).$ (3.13)
$H_{F}$ has the eigenstates
@@
where $n=-N^{2},\cdots,0,\cdots,N^{2}$ specifies the energy level within the Fermi sea. The lowest level of energy is $-NM_{P}/2$ which corresponds to an empty Fermi sea, while the highest energy $NM_{P}/2$ corresponds to an completely filled Fermi sea. In the leading order of large $N$ limit, the fuzzy sphere has the radius
$R=Nl_{P}$ (3.17)
and the total Hamiltonian $H$ can be written as
$H=\frac{M_{P}}{2N}\sum_{a=1}^{N^{2}}\xi_{a}^{\dagger}\xi_{a}+\chi_{a}^{\dagger}\chi_{a},$ (3.18)

### v3

@@ Abstract
We propose a quantum mechanical theory of quantum spaces described by large $N$ noncommutative geometry as a model for quantum gravity. The model admits fuzzy sphere as static solution. Over the fuzzy geometry, the quantum mechanics of the fermions is given by a sum of oscillators with equal frequency. The energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy. This state of the fuzzy sphere obeys the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere is identified with the black hole horizon. Moreover the set of states in the Fermi sea gives precisely the Bekenstein-Hawking entropy. We thus propose that quantum black holes are described by fuzzy spheres with a half-filled Fermi sea in our model. We also consider a system of two fuzzy spheres by embedding them as blocks in the matrix quantum mechanics. When the distance $r$ between the two fuzzy spheres is small, the total energy of the system can be computed using perturbation theory. We show that in the leading order of large $N$ limit, the interaction energy depends on $GM_{1}M_{2}$ exactly the manner as in Newton gravity. To reproduce the correct $r$ dependence in the long range, we expect the inclusion of large $N$ corrections and quantum effects will be needed.
@@ 1 Introduction
where $h(X)$ denotes some Yukawa coupling and $V(X)$ the self-interaction. Our goal was to identify basic properties of quantum gravity which are essential to the construction of the theory. Our hope was that the model would be helpful in ways similar to that of the Bohr atomic model to the building of quantum mechanics. We found [19] that if our model admits a constant density of energy eigenstates and if the Fermi sea of the system is filled up to a Fermi energy level that is inversely proportional to the system size, then the Schwarzschild radius of black hole is reproduced for the system. Moreover the system is in a highly degenerate energy state whose counting of microstates gives precisely the Bekenstein-Hawking entropy. While the result is quite encouraging, the identified conditions are sufficient ones, and it is not clear if all the assumptions can indeed be implemented consistently in a quantum mechanics. The construction of a consistent quantum mechanical model of spacetime which includes quantum black hole is the main motivation of this work.
@@
In this work, we propose an explicit large $N$ quantum mechanics of non-abelian bosonic and fermionic variables as a model of quantum gravity. The use of large $N$ is inspired by the previous success of the BFSS and IKKT matrix models, see, for example, [21, 22] for review. There are three bosonic variables $X^{a}$ in the adjoint of $SU(N)$ which correspond to the three spatial coordinates of a quantized noncommutative space. We note that the emergence of noncommutative geometry is generally expected when space (or spacetime) is quantized in quantum gravity  [^1] The studied of quantized spacetime in terms of noncommutative geometry goes back to the study of Snyder [23] and Yang [24]. . In addition, there are two fermionic variables which couple minimally to the geometry, and make up the fundamental Hilbert space of quantum gravity. We show that our model of quantized spacetime admits fuzzy sphere as static solution. Over the fuzzy geometry, the quantum mechanics of the fermions gives a Fermi sea with uniform energy levels in the large $N$ limit. It is amazing that the energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy and satisfies the the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere is identified with the black hole horizon. Moreover the set of states in the half-filled Fermi sea gives precisely the Bekenstein-Hawking entropy. We therefore propose that quantum black holes are described by fuzzy spheres with a half-filled Fermi sea in our model. Quite amazingly, the quantum mechanics proposed here realizes pretty much all the ideas outlined earlier in [18, 19, 20].
@@
To show that the fuzzy spheres are gravitating objects, one needs to show that the interaction between them agrees with gravity, e.g. with Newton gravity in the long distance limit. We thus consider a system of two fuzzy spheres by embedding them as blocks in the matrix quantum mechanics. When the distance $r$ between the two fuzzy sphere is small, the total energy of the system can be computed using perturbation theory. We show that in the leading order of large $N$ limit, the interaction energy depends on the product of black holes masses and the Newton constant exactly as is expected in Newton gravity. it is possible that the inclusion of large $N$ corrections as well as quantum loop effects could reproduce the correct Newton limit at large distance. We leave this important problem for further study.
@@
The plan of the paper is as follows. In section 2, we give our proposal of a large $N$ quantum mechanics as a fundamental formulation of quantum gravity in 3-dimensions. In section 3, we show that the fuzzy sphere solution with a half-filled Fermi sea has the desired properties of a quantum black hole, namely with the black hole mass-radius relation and the Bekenstein-Hawking black hole entropy reproduced correctly.In section 4, we discuss the mode stability of our fuzzy sphere solution and show that it is stable with respect to mass preserving perturbations. In section 5, we consider a system of 2 fuzzy spheres and derive their interaction energy in the small separation limit. We discuss how the inclusion of large $N$ corrections as well as quantum loop effects could generate the desired gravitational forces between the quantum black holes. Further discussion is found in section 6.
@@ 2 A proposed model of quantum gravity
Let us consider a model of quantized space where the spatial coordinates becomes operators $X^{a}$ ($a=1,2,3$) and are represented by $SU(N)$ matrices. Here $N$ is large and characterizes the dimensions of the Hilbert space of quantum gravity. In addition, we propose to include fermionic degrees of freedom in the fundamental formulation of quantum gravity. In our model proposed below, they are given by the 2-components spinors $\psi^{\dagger},\psi$ in the adjoint representation of $SU(N)$. Our choice of degrees of freedom is a minimalistic one and is allowed since we do not assume supersymmetry. In fact, it is important to realize that going bottom-up, there is no physical reason to insist on having supersymmetry. We remark that in the usual top down approach, supersymmetry is often needed for the reason of consistency of formulation. e.g. for string/M theory in 10 or 11 dimensions. However, supersymmetry is irrelevant for the well-definedness and consistency of quantum mechanics as there is no issue of renormalizability here. Also the usual advantages of supersymmetric quantum mechanics such as exact computation of quantities such as index or partition function does not appeal to us since here we are after the construction of a physical theory rather than a theory that is exact or calculable. Therefore we will not insist on having supersymmetry. This allows for a much wider choice of terms in the Lagrangian which are forbidden otherwise.
We propose to consider the following quantum mechanical model of spacetime, with the Lagrangian
$L=\mbox{tr}\left[\frac{1}{2M_{0}}\dot{X}^{a2}+\frac{M_{P}}{N^{2}}\left([X^{a},X^{b}]^{2}+4X^{a2}\right)+i\dot{\psi}^{\dagger}\psi-a_{2}\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi\right]-a_{3}r_{X}M_{P}$ (2.1)
and the Hamiltonian
$H=\mbox{tr}\left[\frac{M_{0}}{2}P^{a2}-\frac{M_{P}}{N^{2}}\left([X^{a},X^{b}]^{2}+4X^{a2}\right)+a_{2}\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi\right]+a_{3}r_{X}M_{P}.$ (2.2)
where $P^{a}=\dot{X}^{a}/M_{0}$ is the conjugate momentum for the bosonic degrees of freedom. The matrices $X^{a},\psi,\psi^{\dagger}$ are dimensionless $N\times N$ traceless Hermitian matrices, meaning that the model (2.1) is supposed to describe the dynamics with respect to the center of mass frame. $\sigma^{a}$ are the Pauli matrices. We note that unlike the IKKT or BFSS matrix model where the maximal supersymmetry fixes the field content and the action uniquely, our model is nonsupersymmetric and there is a negative mass term. In (2.1), the quartic term is fixed to have a coefficient 4 times that of the mass term. This can always be achieved with a rescaling of the variables $X^{a}$. Here $M_{0}=a_{0}^{2}M_{P}$ and the Lagrangian is specified by a mass scale $M_{P}$ (Planck mass) that sits in front of the bosonic potential, whose relation with the Newton constant $G$ will be specified below. The field $\psi$ is normalized so that the fermionic kinetic term has a unit coefficient. We have adopted a definite $N^{2}$ dependence in the bosonic potential and the Yukawa coupling term. The bosonic kinetic term and the Yukawa term are then specified up to a choice of the coefficients $a_{0}$ and $a_{2}$. In addition, we have included a topological action term specified by $a_{3}$. Here $r_{X}$ is the rank of the matrix $\Gamma:=[X^{a},X^{b}]^{2}$. The $r_{X}$ term does not affect the equation of motion, but measures the energy of space due to noncommutativity. It is $r_{X}=0$ for Abelian configurations (including the Minkowski vacuum $X^{a}=0$) and $r_{X}=N$ for the fuzzy sphere solution. We note that the coefficients $a_{0},a_{2},a_{3}$ are dimensionless. They are not fixed by symmetry argument, but will be determined below “phenomenologically” by requiring the properties of the fuzzy sphere solutions match up with that expected of the quantum Schwarzschild black holes. In the large $N$ limit, our model is equivalent to some non-supersymmetric version of the IKKT-like instantonic model in 4 dimensions by a large $N$ reduction [25].
The Lagrangian (2.1) has an $SO(3)$ rotational invariance where $X^{i}$ transforms as a vector and $\psi$ in the spinor representation. The theory is also invariant under the global $SU(N)$ transformation
@@
We remark that in the SUSY BFSS matrix model, the $SU(N)$ symmetry is usually gauged in order to close the SUSY algebra. In [26], it was shown that the global BFSS matrix QM, although it contains extra non-singlet states that do not come in supersymmetry multiplets, is nevertheless also consistent. In our case, it is also possible to gauge the $SU(N)$ and impose a singlet Gauss law constraint. We do not consider this possibility here but simply noting that the fuzzy sphere solution is also a solution of the gauged model and our analysis below remains the same. We remark that our proposal is similar in philosophy to that of the M(atrix) theory [21]. However we do not assume supersymmetry, and so we can write down our model directly for 3 space dimensions, and we can include a mass term. We note that a mass term has been considered in the IKKT matrix model where interesting classical solutions such as fuzzy spacetime [27] and expanding universe [28, 29, 30, 31], have been obtained. However, the possible connection of a nonsupersymmetric large $N$ quantum mechanics with black hole is new.
3 Fuzzy sphere solution with Fermi sea
@@
$\frac{1}{M_{0}}\ddot{X}^{a}+\frac{4M_{P}}{N^{2}}\left([X^{b},[X^{a},X^{b}]]+2X^{a}\right)=0,\quad\psi=0.$ (3.1)
@@
$X^{a},X^{b}],X^{b}]=2X^{a}.$ (3.2)
This can be solved by the spin j=(N1)/2j=(N-1)/2 representation of $SU(2)$, which are given by $2j+1=N$ dimensional matrices satisfying
$X^{a},X^{b}]=i\epsilon_{abc}X^{c}.$ (3.3)
@@
the configuration (3.3), (3.4) define a fuzzy sphere. Since $r_{X}=N$, the classical fuzzy sphere solution carries an energy
$E_{B}=(2a_{3}-1)\frac{M_{P}N}{2}$ (3.5)
in the leading order limit of large $N$.
Next, let us analysis the Yukawa coupling term
$H_{F}=a_{2}\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi.$ (3.6)
Over the fuzzy sphere geometry, the matrix $K:=\sigma^{a}X^{a}$ satisfies $K^{2}+K-\frac{N^{2}-1}{4}{\bf 1}=0$ and so it has eigenvalues (N1)/2(N-1)/2 or (N+1)/2-(N+1)/2. Since it is traceless, therefore $N+1$ of the eigenvalues are positive and N1N-1 of them are the negative ones. Let us from now on consider the leading large $N$ limit. In the leading order of large $N$ limit, $K$ has the eigendecomposition
@@
Here ${\cal U}^{p}_{n\beta},{\cal V}^{p}_{n\beta},p=1,\cdots,N$ are eigenvectors of $K$ with positive and negative eigenvalue. Introduce the fermionic oscillators
$\xi^{p}_{k}:={\cal U}^{p\dagger}_{n\beta}\psi_{nk\beta},\quad\chi^{p\dagger}_{k}:={\cal V}^{p\dagger}_{n\beta}\psi_{nk\beta}$ (3.8)
and $H_{F}$ reads
$H_{F}=\frac{a_{2}M_{P}}{2N}\sum_{p,k=1}^{N}\left(\xi_{k}^{p\dagger}\xi_{k}^{p}-\chi_{k}^{p}\chi_{k}^{p\dagger}\right).$ (3.9)
Here the Grassmanian variables $\xi_{k}^{p}$ and $\chi_{k}^{p}$ resemble the oscillators $b$ and $d$ of the Dirac theory of electrons.
The theory (2.2) can be canonically quantized. For the bosonic variables, we impose the commutation relation
$X^{a}_{mk},P^{b}_{nl}]=i\delta^{ab}\delta_{mn}\delta_{kl}.$ (3.10)
As for the fermionic variables, the conjugate momentum $\pi=i\psi^{\dagger}$ give rises to the anti-commutation relation
$\{\psi^{\dagger}_{mk\alpha},\psi_{nl\beta}\}=\delta_{mn}\delta_{kl}\delta_{\alpha\beta}.$ (3.11)
This gives equivalently,
$\{\xi^{p}_{k},\xi^{q\dagger}_{l}\}=\delta^{pq}\delta_{kl},\quad\{\chi^{p}_{k},\chi^{q\dagger}_{l}\}=\delta^{pq}\delta_{kl}.$ (3.12)
The quantized fermion Hamiltonian can be obtained from (3.9) with the subtraction of a constant,
$H_{F}=\frac{a_{2}M_{P}}{2N}\sum_{p,k=1}^{N}\left(\xi_{k}^{p\dagger}\xi_{k}^{p}+\chi_{k}^{p\dagger}\chi_{k}^{p}\right)$ (3.13)
such that there is no zero point energy for the oscillators. This may also be obtained with a prescription of normal ordering defined on the oscillators $\xi^{p}_{k}$ and $\chi^{p}_{k}$, $\chi^{p}_{k}\chi^{p\dagger}_{k}:=-\chi^{p\dagger}_{k}\chi^{p}_{k}$ etc. The quantized Hamiltonian $H_{F}$ has the eigenstates
@@
where $n=-N^{2},\cdots,N^{2}$ specifies the energy level within the Fermi sea. The lowest level $n=-N^{2}$ corresponds to an empty Fermi sea, while the highest level $n=N^{2}$ corresponds to an completely filled Fermi sea. We remark that one may label the oscillators with the collective indices $a:=(p,k)$ and the fermionic Hamiltonian can be written as
$H_{F}=\frac{a_{2}M_{P}}{2N}\sum_{a=1}^{N^{2}}\xi_{a}^{\dagger}\xi_{a}+\chi_{a}^{\dagger}\chi_{a}.$ (3.17)